OpenStudy (anonymous):

are f(x)=2x-4 and g(x)=1/2x+2 inverse of each other?

OpenStudy (campbell_st):

to find the inverse of y = 2x - 4 swap x and y x = 2y - 4 now make y the subject

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=2x-4 y+4=2x x=(y+4)/2 interchange x and y: y=(x+4)/2